ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 16 Moussavi, Ahmad, University of Tarbiat Modares, Iran On The Semiprimitivity of R[X; F, D] Let R be a ring, F : R → R be a monomorphism and D : R → R be a Dderivation. We show that if R is semiprime left Noetherian, then R[X; F, D] is semiprimitive. Reference: Moussavi, A. On the semiprimitivity of skew polynomial rings, Proc. of the Edinburgh math. Soc. (1993) 36, 169-178.